### v1

# Unmasking Deepfakes: Leveraging Augmentations and Features Variability for Deepfake Speech Detection

Abstract
The detection of deepfake speech has become increasingly challenging with the rapid evolution of deepfake technologies. In this paper, we propose a hybrid architecture for deepfake speech detection, combining a self-supervised learning framework for feature extraction with a classifier head to form an end-to-end model. Our approach incorporates both audio-level and feature-level augmentation techniques. Specifically, we introduce and analyze various masking strategies for augmenting raw audio spectrograms and for enhancing feature representations during training. We incorporate compression augmentations during the pretraining phase of the feature extractor to address the limitations of small, single-language datasets. We evaluate the model on the ASVSpoof5 (ASVSpoof 2024) challenge, achieving state-of-the-art results in Track 1 under closed conditions with an Equal Error Rate of 4.37%. By employing different pretrained feature extractors, the model achieves an enhanced EER of 3.39%. Our model demonstrates robust performance against unseen deepfake attacks and exhibits strong generalization across different codecs.
keywords:
Deepfake Speech Detection , Speech Processing , ASVSpoof5 , Speech Augmentations , Speech Features
journal: Speech Communication
1 Introduction
1.1 Deepfake Speech Detection
Deepfake speech poses serious security concerns across various fields, including cybersecurity, law enforcement, and military operations. Synthetic audio can be exploited for misinformation, impersonation, and fraud, necessitating robust detection techniques to ensure the integrity of audio-based communications. Initial detection methods centered on traditional feature extraction techniques, such as Mel-frequency cepstral coefficients (MFCC) and spectrogram analysis, which capture essential acoustic properties to help differentiate genuine from synthetic speech [1, 2, 3, 4]. These traditional features were often combined with classifiers like support vector machines (SVMs) and Gaussian mixture models (GMMs), achieving initial successes in synthetic speech detection through the modeling of fundamental spectral patterns [5, 6]. Deep learning approaches, including convolutional neural networks (CNNs) and recurrent neural networks (RNNs), have advanced deepfake speech detection. CNNs, particularly ResNet models, capture local acoustic features [4, 7], while RNNs like LSTMs model temporal dependencies for improved classification [8]. Hybrid CNN-RNN architectures combine spatial and sequential features effectively [9].
The ASVSpoof challenge series [10, 11] has become a critical benchmark for developing and evaluating anti-spoofing techniques in Automatic Speaker Verification (ASV) systems, adapting over time to address increasingly sophisticated deepfake threats. The latest iteration, ASVSpoof5, consists of two tracks: Track 1 focuses on standalone deepfake speech detection, independent of ASV systems, while Track 2 integrates spoof detection within ASV pipelines to mirror real-world applications. ASVSpoof5 includes diverse and challenging datasets with various synthetic methods, codecs, and speaker variations, evaluated by Equal Error Rate (EER) and additional metrics to measure detection accuracy. By continuously raising the standard for spoof detection, ASVSpoof encourages advancements in robust, generalizable solutions to counter deepfakes, which are critical in high-stakes fields such as cybersecurity and law enforcement [12].
1.2 Speech Features
Audio features play a critical role in deepfake speech detection by providing a more structured representation of the raw audio signal, which is inherently a one-dimensional vector. Extracting meaningful features from audio, such as spectral, temporal, or statistical properties, enables models to better capture patterns and anomalies associated with deepfake generation processes
Spectrograms are a classic feature extraction technique that provide a time-frequency representation of the audio signal. By displaying the amplitude of frequencies over time, spectrograms enable models to capture not only the phonetic content but also fine-grained details of the audio, including artifacts introduced by deepfake generation methods such as speech synthesis and voice conversion [13]. Spectrograms have proven effective in deepfake detection, as mentioned in [14] and [15], CNN-based architectures applied to spectrogram inputs can significantly differentiate between genuine and spoofed speech. However, despite their success, the use of spectrograms may be approaching a performance ceiling. Recent advances suggest that further improvements in deepfake detection may require more sophisticated feature extraction methods, as the complexity of spoofing attacks continues to increase.
Wav2Vec, a deep learning-based feature extractor introduced in [16], was originally designed for automatic speech recognition (ASR) tasks. Wav2Vec 2.0, a refined version [17], leverages self-supervised learning to generate high-level speech representations directly from raw audio. Although initially developed for ASR, recent research has demonstrated the efficacy of Wav2Vec features in deepfake detection tasks [18, 19]. By utilizing self-supervised learning, Wav2Vec captures rich contextual information from speech, which is essential for identifying subtle anomalies in deepfake speech that traditional methods may miss. These representations, being task-agnostic and capable of generalizing across diverse conditions, have shown promise in enhancing deepfake detection systems against both known and novel attacks.
1.3 Speech Augmentations
Data augmentation techniques play a pivotal role in enhancing the robustness of deepfake speech detection models. By introducing variations to the training data, augmentation methods aim to simulate real-world conditions, enabling the models to learn generalized patterns that can better handle unseen, adversarial, or out-of-distribution inputs. Common speech augmentation techniques include pitch shifting, time-stretching, noise addition, and speed perturbation. These methods help models adapt to different speaker characteristics, environmental conditions, and recording qualities [3, 20].
Beyond speech augmenting, additional methods focus on audio augmenting. SpecAugment [21], developed for speech recognition, enhances generalization by masking random time and frequency blocks in the spectrogram. A further adaptation, SpecAverage [7], was proposed to address the non-zero mean characteristics of audio features, making it more suitable for deepfake speech detection. However, a limitation of both SpecAugment and SpecAverage is that the applied masks have a fixed shape, which may not adequately represent the complex and dynamic noise distributions encountered in real-world speech data. This constraint arises from the fact that the mask shapes are primarily designed for computational efficiency rather than mimic audio distortion patterns.
1.4 Main Contributions
This work presents a robust model for deepfake speech detection, trained end-to-end on a limited dataset from the ASVSpoof5 dataset, achieving SoTA results in Track 1 under the closed conditions. We specifically focus on data augmentation and training techniques that enable the development of a resilient model despite the limitations of the dataset. The key contributions of this paper are as follows:
  1. 1.
    We introduce novel augmentation strategies applied at multiple stages of the audio processing pipeline. We propose the use of raw audio augmentations, which are first applied to the raw audio signal, converted to a spectrogram, and then processed as input to the model. Furthermore, we extend this approach by introducing feature-level augmentations that are applied directly to the output of the feature extractor, in the middle of the pipeline.
  2. 2.
    We integrate audio compression augmentations into the pretraining phase of the feature extractor to strengthen its robustness. This approach is particularly important in the context of self-supervised learning, especially when working with small, single-language datasets. By exposing the model to compression artifacts during pretraining, we introduce variability into the input data, which is critical for self-supervised learning. This enhances the model’s ability to generalize and improves its adaptability to new, unseen conditions.
  3. 3.
    We propose a hybrid architecture that combines Wav2Vec 2.0 structure as feature extraction with a ResNet structure as classification head. This architecture is trained using the power of self-supervised learning for feature extraction pertaining and supervised learning for training the model end-to-end for deepfake speech detection.
2 Method
This section outlines the methodology we employed, beginning with our proposed hybrid architecture, which combines a trainable feature extraction module (Wav2Vec2) with a ResNet34 classification head. Our approach incorporates augmentation strategies applied at different stages of the pipeline to enhance model performance. For raw audio, we employ MaskedSpec, low-pass filtering (LPF), and compression augmentation. Compression augmentations, in particular, are utilized both during the pretraining phase of the feature extractor and in the end-to-end training process. At the feature level, our work introduces MaskedFeature and feature normalization.
Figure 1: Illustration of our method: raw audio serves as the input, followed by raw-audio augmentations. The augmented audio is processed through a feature extractor, after which the extracted features undergo additional augmentations. Finally, the classification head processes the enhanced features to generate predictions.
2.1 Hybrid model
We propose a hybrid model that combines a trainable feature extractor with a classification head, designed to achieve joint optimization of feature extraction and classification. The training process comprises two phases: (1) a self-supervised pretraining phase for the feature extractor, enabling it to learn robust and transferable latent representations from raw audio data, and (2) an end-to-end supervised training phase for the entire hybrid model, ensuring that the extracted features align with task-specific classification objectives. As illustrated in Figure 1, augmentations are applied at multiple stages during the end-to-end training process, targeting both raw audio inputs and extracted features. The hybrid design integrates two sub-models, with each sub-model benefiting from augmentation strategies.
Leveraging the success of Wav2Vec 2.0 in speech recognition tasks, we adopted it as the feature extractor in our architecture. This choice is motivated by Wav2Vec 2.0’s ability to capture meaningful audio representations without the need for extensive labeled data, thanks to its self-supervised learning paradigm. For the classification head, we employed ResNet34, a deep convolutional neural network renowned for its robust performance in image classification and its proven adaptability for audio classification tasks, including deepfake detection using traditional frequency-domain features.
2.2 Augmentation Strategies
To address the challenges posed by unseen spoof attacks and varying codecs, we suggest augmentation strategies that aim to introduce variability during training, enabling the model to generalize better and adapt to diverse conditions  [^1] https://github.com/InbalRim/MaskedSpec. Our methodology incorporates augmentations at two distinct stages of the processing pipeline to address different aspects of corruption. First, augmentations are applied directly to the raw audio signals, modifying the dataset itself. By introducing such corruptions at the raw data level, the model is exposed to a broader range of input variations, improving its resilience. Second, augmentations are applied to the classifier input, focusing on the audio features and intermediate representations derived from the raw data. These feature-level augmentations simulate distortions that occur in processed audio, helping the model learn to recognize and handle nuanced variations in audio representations.
2.2.1 Raw Audio Augmentations
Figure 2: Visualization of different mask types used in the experiments: Squares, Bands, Singles, and Gauss.
Inspired by SpecAugment [21] and SpecAverage [7], we sought to enhance the training dataset by performing augmentations on the raw audio through its frequency domain representation. The proposed MaskedSpec technique employs frequency and time-domain masking to the audio’s Short-Time Fourier Transform (STFT). First, the STFT of the raw audio signal is calculated using the formula:
$X(m,k)=\sum_{n=0}^{N-1}x(n+m)\cdot w(n)\cdot e^{-j2\pi kn/N},$
where $x(n)$ is the input audio, $w(n-m)$ represents the window function, $m$ and $k$ are the time and frequency bins, and $N$ is the frame size.
Next, the mean value of the STFT is computed. Initially, the mean magnitude and mean phase of the STFT are calculated as follows:
$\mu_{\text{magnitude}}=\frac{1}{N}\sum_{n=1}^{N}|x(n)|$ (1)
$\mu_{\text{phase}}=\frac{1}{N}\sum_{n=1}^{N}\arg(x(n))$ (2)
Then, a complex value is reconstructed from these means:
$\mu_{\text{stft}}=\mu_{\text{magnitude}}\cdot\exp(j\cdot\mu_{\text{phase}})$ (3)
$\mu_{\text{stft}}$ is a complex number that serves as the masking value to occlude portions of the STFT. After the masks are applied, the inverse Short-Time Fourier Transform (iSTFT) is used to convert the modified STFT back into the time domain, and the resulting signal is then passed as input into the feature extractor for further processing.
Several masking shapes were explored to optimize the model’s capacity for generalization: Square masks occlude localized regions of the spectrogram, encouraging reliance on global patterns over specific, localized features. Bands masks introduce elongated occlusions across continuous frequency ranges, mimicking spectral loss and requiring the model to capture dependencies over extended temporal or spectral contexts. Singles masks target individual frequency bands, providing simpler occlusions that still effectively challenge the model. Gaussian mask gradually degrade the spectrogram, simulating natural signal distortions such as noise-induced corruption or environmental degradation. While traditional masking strategies are often favored for their computational efficiency and ease of implementation, we sought to explore masks that, although more computationally demanding, could offer improved relevance and effectiveness for the task. These varied masking strategies aim to expose the model to a wide range of distortions, improving its robustness to unseen conditions. A visualization of these masking types is presented in Figure 2.
Low-pass filtering (LPF) is a fundamental signal processing technique often employed to simulate the loss of high-frequency information in various audio applications. In the context of deepfake speech detection, LPF is particularly effective in modeling the compression or recording artifacts that frequently occur in real-world scenarios. By attenuating the higher-frequency components of the audio signal, LPF mimics the effects of common distortions, such as those introduced by lossy audio codecs or imperfect recording conditions. This augmentation method plays a crucial role in improving the robustness and generalization ability of the model. Additionally, the randomization of cutoff frequencies used in LPF further diversifies the training data, exposing the model to a broader range of possible high-frequency losses. This, in turn, enhances the model’s ability to generalize across different audio conditions. We perform LPF using 1D convolution operation applied to the audio signal $x[n]$ with a filter kernel $h[n]$, which acts as the low-pass filter. The filtered output signal $y[n]$ is given by the following convolution formula:
$y[n]=\sum_{m=0}^{M-1}x[n-m]h[m]$ (4)
The filter kernel, $h[m]$, is constructed using a sine function modulated by the cutoff frequency, $f_{\text{cutoff}}$, and smoothed with a Hamming window, $\text{window}[m]$, to minimize spectral leakage. The formula for $h[m]$ is given by:
$h[m]=\frac{\sin(2\pi f_{\text{cutoff}}m)}{\pi m}\cdot\text{window}[m],$
where a special case is applied at $m=0$ to prevent division by zero:
$h[0]=2\cdot f_{\text{cutoff}}.$
The use of the Hamming window reduces edge discontinuities, ensuring a smooth and stable frequency response. This design effectively attenuates frequencies above the cutoff, preserving the desired range and improving overall filter performance.
In order to address the challenges posed by real-world audio distortions, particularly from commonly used lossy compression, we incorporated a compression-decompression (codec) augmentation into our pipeline. Lossy compression algorithms such as MP3 and M4A are widely used for audio storage and streaming, due to their ability to reduce file sizes by discarding some of the less perceptible audio information. However, the compression process often introduces artifacts such as quantization noise, distortion, and loss of high-frequency content, which can negatively impact audio quality. To simulate the effects of these distortions, we applied the augmentation by encoding and decoding the raw audio at various bitrates. This encoding process generates artifacts unique to the compression algorithm used, such as the introduction of audible noise and spectral smearing, which occur due to the limited bitrate and reduced signal fidelity. In our experiments, we used the compression augmentation in both the end-to-end supervised training step and the pretraining self-supervised phase. In the end-to-end training, the model is exposed to audio samples subjected to compression and decompression at varying bitrates, compelling it to learn to detect deepfake speech even when compression artifacts are present. For self-supervised pretraining, training relies on a diverse and extensive dataset to ensure that the model can learn meaningful representations from a broad range of input variations. The variability introduced by compression artifacts further enriches the dataset, providing additional diversity and improving the model’s ability to generalize. By exposing the model to a wider variety of noise and distortion types, the pretraining process is strengthened, leading to better performance when the model is later fine-tuned for specific tasks, such as deepfake detection.
2.2.2 Feature Augmentations
To further expose the model to potential feature artifacts and improve its generalization ability, we employed an augmentation strategy similar to MaskedSpec, but applied directly to the latent representations produced by the feature extractor. This augmentation is referred to as MaskedFeature. The primary goal of this technique is to force the model to rely on broader, more generalized patterns in the data, rather than focusing on specific, localized details that might be overly sensitive to noise or distortions. In MaskedFeature, portions of the feature space are occluded during training, which simulates the presence of distortions or information loss in the latent representations. By masking these parts of the feature space, the model learns to adapt to a more holistic view of the input data, promoting robustness and resilience. This approach is particularly beneficial in scenarios where certain features of the input may be corrupted or unavailable, such as in deepfake detection tasks where some parts of the signal may be intentionally manipulated. We also investigate the use of various mask shapes for MaskedFeature, aiming to further explore how different patterns of occlusion influence the model’s ability to generalize.
Feature normalization was implemented to scale the extracted feature values into a consistent range, such as $[-1,1]$. Normalization minimizes the dominance of individual features, promoting balanced learning and reducing sensitivity to recording inconsistencies or device-specific biases. This step is critical for ensuring robustness across varied input conditions.
3 Results and Insights
This section presents the experimental results of the proposed methods, beginning with an overview of the experimental setup and evaluation metrics. We then evaluate the architecture’s performance and training process while experimenting with various augmentation techniques, focusing on examining the role of MaskedSpec and MaskedFeature augmentations in improving model robustness and generalization across diverse conditions. Next, we examine the role of pretraining the feature extractor in improving representation quality. Finally, we performed a fusion of the top-performing models, each trained with distinct augmentation strategies and pretraining configurations.Models trained with MaskedSpec using the Bands mask and MaskedFeature with the Gauss mask, alongside different pretraining weights for the feature extractor, were fused to achieve an EER of 3.39%. This result highlights the value of integrating our independent augmentation strategies that capture diverse perspectives on the data during training, thereby enhancing the model’s robustness and overall performance.
3.1 Experiments Setup
Experiments were conducted using the proposed architecture model, exploring the effects of various data augmentation and manipulation techniques. Models trained on a GPU with 24 GB of memory. The training utilized a constant batch size of 16, and the number of epochs was limited to 5 to prevent overfitting while addressing time constraints. Compression augmentations employed MP3 with a bitrate of 16 kbps and M4A at 64 kbps during training. For development, we used MP3 at 48 kbps along with M4A at both 16 kbps and 64 kbps. Although development augmentations are typically not performed, their inclusion in this case enhances the assessment of the model’s robustness. LPF was applied online, with a randomly selected frequency cutoff. For both MaskedSpec and MaskedFeature, masks were generated online, with a random selection of the number and size of each patch. Although creating masks online is time-consuming, this approach is valuable for capturing diverse aspects of data manipulation.Pretrained feature extractors were used solely for initialization, with the entire model, including the feature extractor, subsequently trained end-to-end throughout the training process. The evaluation is conducted using the ASVSpoof5 datasets, where the Equal Error Rate (EER) for both the development (Dev EER) and evaluation (Eval EER) sets is calculated. Although other metrics within the ASVSpoof5 framework provide useful insights, we prioritize EER due to its significance in assessing the model’s ability to accurately differentiate between genuine and deepfake speech.
3.2 Model Architecture and Features
Features Dev EER (%) Eval EER (%)
Spectrogram 47.86 38.72
Wav2Vec2 (without pretraining) 38.24 40.64
Wav2Vec2 (pretrained on ASVSpoof5) 20.82 26.36
Table 1: Performance of classification head with different features.
We explore the performance differences between traditional, non-trainable features such as spectrograms and trainable feature extractors like Wav2Vec2, using the same classification head, ResNet34. Traditional features like spectrograms provide a static representation of audio based on predefined transformations, requiring the classifier to identify meaningful patterns from these fixed inputs. On the other hand, trainable feature extractors, such as Wav2Vec2, dynamically learn task-specific representations directly from raw audio, leveraging deep learning’s ability to extract hierarchical features.
The results in Table 1 demonstrate the clear advantages of trainable features over traditional spectrogram-based representations. Models trained with spectrogram features exhibit the weakest performance, with a Dev EER of 47.86% and Eval EER of 38.72%. This indicates that spectrograms, without additional learned enhancements, struggle to effectively separate genuine from spoofed audio in this task. Introducing Wav2Vec2 as the feature extractor significantly reduces the EER, even without pretraining, achieving a Dev EER of 38.24% and Eval EER of 40.64%. Despite this improvement, the lack of domain-specific pretraining likely limits its generalization capabilities.
Significant improvement is observed when the model is trained in two distinct phases. Initially, Wav2Vec2 undergoes self-supervised pertaining, as detailed in [22], on the ASVSpoof5 dataset, allowing it to learn meaningful representations of audio data without relying on labels. Subsequently, end-to-end training is performed on the hybrid model using labeled data, fine-tuning both the feature extractor and the classification head for the specific task. This dual-phase training approach achieves a notable reduction in error rates, resulting in a Dev EER of 20.82% and an Eval EER of 26.36%, showcasing the combined benefits of unsupervised pretraining and supervised fine-tuning.
3.3 MaskedSpec and MaskedFeature Experiments
Mask Type Dev EER (%) Eval EER (%)
None 2.86 6.15
Squares 1.73 6.80
Bands 1.36 4.53
Singles 1.56 5.58
Gauss 3.89 6.06
Table 2: Model performance comparison of different MaskedSpec masking strategies.
In the following experiment, we evaluate the model’s performance when trained with either MaskedSpec or MaskedFeature augmentation techniques. These experiments are conducted using the XLS-R 53 pertaining for the feature extractor [23], in combination with compression augmentations. The aim is to assess how the different masking strategies affect the model’s ability to generalize and handle artifacts introduced by compression, ultimately improving its performance in detecting deepfake speech.
The results summarized in Table 2 highlight the effectiveness of MaskedSpec augmentation with various masking strategies on the model’s performance, compared to a baseline model trained without any MaskedSpec augmentation. The Bands masking strategy showed best performance, achieving the lowest EERs on both the development set (1.36%) and the evaluation set (4.53%), outperforming the baseline in both metrics.
While the Squares mask demonstrated a competitive development EER of 1.73%, its evaluation EER of 6.80% indicates less consistency when tested on unseen data, suggesting that its improvements are more confined to the training set conditions. The Singles mask presented a balanced performance, with moderately low EERs on both the development (1.56%) and evaluation (5.58%) sets. In contrast, the Gauss mask exhibited a higher development EER of 3.89% but still slightly improved the evaluation EER to 6.06% compared to the baseline. The improvement in evaluation EER over the baseline indicates that the Gauss mask still contributes positively to the model’s ability to handle real-world distortions, albeit less consistently than strategies such as Bands or Singles. This highlights the trade-off between overfitting to development conditions and generalizing to unseen data, emphasizing the need to balance augmentation techniques for optimal performance across datasets.
MaskedSpec Feature Norm LPF Dev EER (%) Eval EER (%)
Squares 2.5401 7.64
1.5827 6.05
Bands 1.4808 5.51
1.7414 6.75
Singles 2.4847 5.25
1.9191 7.20
Gauss 0.9402 5.82
2.0702 5.94
Table 3: Model performance with mixed augmentations, showcasing the effects of various MaskedSpec masking strategies combined with feature normalization and low-pass filtering (LPF).
We further examined the combined effects of MaskedSpec masking strategies with low-pass filtering (LPF) and feature normalization, as shown in Table 3. The Gauss mask combined with feature normalization achieved the lowest development EER (0.94%), demonstrating improved training performance over MaskedSpec alone. This configuration also showed an improvement in evaluation EER, reducing it from 6.06% without feature normalization to 5.82%. Conversely, while feature normalization enhanced the evaluation EER for the Singles mask, it led to a slight increase in development EER, indicating a trade-off between training and evaluation performance. Both the Squares and Gauss masks paired with LPF achieved reduced development and evaluation EERs, highlighting their adaptability to unseen data. However, not all mask and augmentation combinations were equally effective; for example, the Bands mask did not yield as robust results, underscoring the need for carefully tailored augmentations to optimize performance across both training and evaluation scenarios.
Mask Type Dev EER (%) Eval EER (%)
None 2.86 6.15
Squares 2.4972 6.73
Bands 1.4397 8.78
Singles 2.0733 13.13
Gauss 1.3802 5.22
Table 4: Model performance comparison of different MaskedFeature masking strategies.
In Table 4, we present the performance of different masking strategies within the MaskedFeature augmentation framework compared to the baseline model without masking. All masking strategies outperform the baseline on the development set, demonstrating the effectiveness of MaskedFeature in reducing the Dev EER. For instance, the Gauss mask achieves the lowest Dev EER of 1.38%, followed closely by the Bands mask at 1.44%. Similarly, the Singles mask and Squares mask also show improvements, reducing the Dev EER to 2.07% and 2.50%, respectively.
However, the performance on the evaluation set tells a different story. While the Gauss mask maintains competitive performance with the lowest Eval EER of 5.22%. The Bands mask, despite its strong Dev EER, sees a significant drop in generalization ability with a high Eval EER of 8.78%. Similarly, the Singles mask, which performed moderately well during development, exhibits poor generalization with an Eval EER of 13.13%. The Squares mask strikes a balance with an Eval EER of 6.73%, improving over some masks but still falling short in terms of robust generalization. These results highlight that while MaskedFeature augmentation effectively reduces the Dev EER, its ability to generalize to unseen data varies significantly across masking types. The Gauss mask emerges as the most promising option for balancing development and evaluation performance, indicating a potential for enhancing robustness in practical scenarios.
3.4 Pretraining Feature Extractor
Feature Extractor Pretrainig Compressions $MS_{B}$ $MF_{G}$ Dev EER (%) Eval EER (%)
XLS-R 53 3.20 8.34
2.86 6.15
ASVSpoof5 20.82 26.36
16.03 20.51
XLS-R 53 1.35 4.53
1.38 5.22
XLS-R 128 0.51 5.06
2.37 5.79
Large-960h 6.03 6.05
9.75 8.25
ASVSpoof5 12.49 12.66
17.03 14.17
ASVSpoof5+ 5.33 5.11
5.96 5.25
Table 5: Evaluation of model performance across various pretrained feature extractors and augmentation strategies. The overall best results are highlighted in bold, with the top-performing results for the ASVspoof 5 Track 1 Closed Condition marked in light gray
We conducted end-to-end training of models initialized with various pretrained weights, using the top-performing configurations for MaskedSpec and MaskedFeature augmentations: Bands and Gauss, respectively. These augmentation configurations are referred to as $MS_{B}$ and $MF_{G}$ in the following sections. The experimental results are summarized in Table 5. To establish a baseline for comparison with subsequent experiments, we first evaluate the model’s performance without augmentation or pertaining the feature extractor. This provides a reference point for assessing the impact of various enhancements.
We investigate the effects of pretraining the feature extractor using XLS-R 53 and ASVSpoof5, both with and without compression augmentations. XLS-R 53 utilizes publicly available weights trained on a diverse dataset, including Multilingual LibriSpeech, CommonVoice, and Babel, comprising over 50k hours of audio. We pretrained a feature extraction model using the ASVSpoof5 dataset, which contains bona fide and spoofed audio in English only and is significantly smaller in scale compared to the datasets used for XLS-R 53. As expected, XLS-R 53 pretrining substantially outperforms ASVSpoof5 pretrining due to its larger and more diverse data. Furthermore, the inclusion of compression augmentations consistently improves performance across all configurations.
Furthermore, we trained the model using the $MS_{B}$ and $MF_{G}$ augmentations. We experimented with additional pretrained weights, including XLS-R 128 [24], which was trained on 436k hours of unlabeled speech data across 128 languages, leveraging datasets such as VoxPopuli, MLS, CommonVoice, BABEL, and VoxLingua107. Another feature extractor. Large-960h, was trained on 960 hours of English-only LibriSpeech data. Additionally, we pretrained with ASVSpoof5 extended with codec augmentations, referred to as ASVSpoof5+.
The results underscore the substantial impact of the pretraining and the augmentation on model performance. Using XLS-R 53 weights, training with $MS_{B}$ and $MF_{G}$ augmentations yielded performance improvements of 26.34% and 15.12%, respectively, compared to models trained with compression augmentations only. Even greater improvements were observed with ASVSpoof5 pretraining, where $MS_{B}$ and $MF_{G}$ augmentations enhanced performance by 38.27% and 30.91%, respectively. Notably, incorporating compression augmentations during pretraining further improved performance, achieving significant gains of 75.08% and 74.4% with $MS_{B}$ and $MF_{G}$, respectively.
Incorporating compression augmentations into pretraining demonstrated a notable impact, yielding approximately a 60% improvement in model performance. This enhancement was observed when comparing models pretrained with ASVSpoof5 to those pretrained with ASVSpoof5+ using either $MS_{B}$ or $MF_{G}$ augmentations.
3.5 Fusions
ASVSpoof5+ XLS-R 53 XLS-R 128 Eval EER (%)
$MS_{B}$ $MF_{G}$ $MS_{B}$ $MF_{G}$ $MS_{B}$ $MF_{G}$
4.37
3.61
3.39
Table 6: Fusions results.
The results presented in Table 6 demonstrate the performance of various fusions of the single models presented in Table 5. The best evaluation EER of 3.39% is achieved by combining models where the feature extractors were pretrained on XLS-R 53, XLS-R 128, and ASVSpoof5+, utilizing both $MS_{B}$ and $MF_{G}$ augmentations. Under ASVSpoof5 closed-conditions, the fusion achieved an EER of 4.37% using $MS_{B}$ and $MF_{G}$ augmentations with feature extractor pretrained on ASVSpoof5+, establishing SoTA results for this benchmark.
4 Analysis of Model Performance Across Deepfake Attacks
Our assessment of model performance relied on insights provided by the competition organizers. These insights included trends in EER across different codecs and spoofing attacks, enabling us to gauge model efficacy without direct access to specific labels. Our observations reveal variation in detection difficulty across attack types, with certain spoofing techniques consistently associated with lower EERs, indicating they are more readily identified by most models.
Model A17 A18 A19 A20 A21 A22 A23 A24 A25 A26 A27 A28 A29 A30 A31 A32
XLS-R 128, $MS_{B}$ 1.40 3.86 2.23 2.04 0.93 2.10 3.61 3.87 1.65 1.61 4.79 10.78 0.99 6.95 9.20 3.16
XLS-R 128, $MF_{G}$ 2.37 6.66 3.39 3.59 1.89 2.82 3.79 6.22 2.32 2.77 9.44 12.57 1.78 12.98 12.43 7.14
XLS-R 53, $MS_{B}$ 1.62 2.68 1.32 1.33 1.56 2.34 3.53 6.03 1.42 1.41 1.76 17.32 1.42 3.76 5.10 1.47
XLS-R 53, $MF_{G}$ 2.00 4.53 2.01 1.75 2.22 2.58 5.05 8.71 1.63 1.94 2.09 16.30 1.74 5.52 6.60 1.56
ASVSpoof5+, $MS_{B}$ 1.70 0.72 5.61 5.01 2.03 4.68 1.64 7.80 3.98 1.29 1.10 19.48 0.84 1.27 4.40 1.66
ASVSpoof5+, $MF_{G}$ 2.19 0.85 5.69 5.25 3.30 5.48 3.37 7.67 4.44 1.84 1.00 22.79 1.04 1.46 4.36 1.52
ASVSpoof5+ fusion 1.19 0.57 3.98 3.72 1.87 3.60 1.91 6.61 3.05 1.00 0.74 19.53 0.74 0.90 3.24 1.02
fusion 1.01 1.82 1.36 1.27 1.23 1.60 2.27 3.50 1.10 1.20 1.79 8.77 0.92 2.53 4.86 1.25
Table 7: Pooled EER performance analysis for spoof attacks. The lowest EER for each spoof attack among single models is highlighted in bold, while the second-lowest is marked in light gray.
The deepfake attacks, denoted A17–A32, presented a range of detection challenges. Attacks A17, A19, A20, A21, A25, A26, and A29 generally produced low EERs across models, suggesting either a relative ease of detection or high compatibility of augmentation methods with these attack types. In contrast, attacks A28, A30, and A31 posed significant detection challenges, as indicated by higher EERs across the majority of tested models. In particular, attack A28 yielded notably elevated EERs, highlighting it as one of the most challenging spoofing types. As presented in Table 7, models utilizing XLS-R 53 pretrained feature extractor, trained with $MS_{B}$, demonstrated robust performance across a broad spectrum of spoofing methods. ASVSpoof5+ achieved lowest EERs on A30 and A31. For A28, identified as the most difficult spoofing type, the model trained with XLS-R 128 pretrained feature extractor and $MS_{B}$ achieved the lowest EER, closely followed by the XLS-R 128 trained with $MF_{G}$. These findings underscore the benefits of augmentations on both audio spectrogram and feature levels, as well as the importance of leveraging pretrained feature extractor weights for improved deepfake speech detection.
When analyzing the fusion of all single models, it is evident that this approach significantly improves performance for attacks such as A17, A20, A22, A24-26, A28, and A32, achieving the lowest EER compared to any single model. For other spoofing attacks, the performance of the fusion model remains comparable to the lowest EER achieved by a single model. Interestingly, the ASVSpoof5+ fusion outperforms the full fusion model on A18, A23, A26-27, A29-32, but demonstrates considerably poorer performance on A28.
Model None C1 C10 C11 C2 C3 C4 C5 C6 C7 C8 C9
XLS-R 128, $MS_{B}$ 1.73 3.10 5.57 2.11 2.89 4.29 10.88 1.96 2.05 13.83 5.90 3.76
XLS-R 128, $MF_{G}$ 3.22 6.02 9.94 3.88 7.07 7.31 12.77 3.31 4.09 15.38 7.58 7.75
XLS-R 53, $MS_{B}$ 1.52 3.47 5.88 3.23 2.97 4.39 7.92 1.69 2.65 9.77 6.37 4.38
XLS-R 53, $MF_{G}$ 1.94 3.61 6.80 3.67 3.52 5.05 9.97 2.00 2.62 12.36 6.72 4.97
ASVSpoof5+, $MS_{B}$ 3.60 4.73 6.49 3.95 4.74 5.69 6.41 4.38 4.67 7.28 6.94 5.77
ASVSpoof5+, $MF_{G}$ 3.88 5.47 7.53 4.62 5.44 7.66 7.30 4.25 4.68 7.81 7.51 6.36
ASVSpoof5+ fusion 3.03 4.26 6.15 3.64 4.26 5.69 5.92 3.59 3.90 6.54 6.23 5.28
fusion 0.59 1.62 3.77 1.21 1.59 2.39 7.53 0.74 0.89 9.55 3.70 2.38
Table 8: Pooled EER performance analysis for codecs. The lowest EER for each codec among single models is highlighted in bold, while the second-lowest is marked in light gray.
Examining detection difficulty across codecs revealed that C4 and C7 presented particular challenges, yielding consistently high EERs across models. This suggests that these codecs may mask or distort essential audio features that models rely on. Conversely, C2, C5, C6, and None (uncompressed) audio exhibited lower EERs, indicating that models were generally able to process and identify deepfake more effectively within these formats. This highlights the critical role of codec selection in the design and training of deepfake speech detection models. The results presented in Table 8 reveal that models utilizing $MS_{B}$ augmentations consistently outperform those trained with $MF_{G}$, indicating that $MS_{B}$ contributes to enhanced robustness across different codecs. While XLS-R 53 and XLS-R 128 generally achieve superior performance compared to ASVSpoof5+, the latter demonstrates notable effectiveness on C4 and C7, underscoring its resilience in challenging scenarios.
The fusion of all single models results in a more robust overall performance, significantly improving upon the results of single models for most codecs. However, for the most challenging codecs, C4 and C7, the fusion achieves good results but does not outperform one or more of the single models. The fusion of ASVSpoof5+ models consistently yields better performance than any individual ASVSpoof5+ model, demonstrating that combining models trained with $MS_{B}$ augmentation and those trained with $MF_{G}$ augmentation provides different perspectives on the data, resulting in a strong overall fusion performance.
5 Discussion
5.1 MaskedSpec vs MaskedFeature
Our results insights into the role of feature and spectral augmentations in deepfake speech detection. Feature-based augmentations (MaskedFeature) and raw-audio-based augmentations (MaskedSpec) each play distinct roles in model robustness. MaskedFeature focuses on perturbing the extracted feature space, forcing the model to adapt to variations in learned representations. MaskedSpec affects the raw audio, directly altering the signal’s spectral and temporal characteristics, thereby exposing the model to a broader range of distortions. The comparison between MaskedSpec (Table 2) and MaskedFeature (Table 4) highlights distinct impacts when applying the same masking strategies to different input representations. In MaskedSpec, where masks are applied directly to the audio input, Bands mask show lowest Eval EER values of 4.53%. Conversely, MaskedFeature masking on the feature-level representations shows a notable shift: the Gauss mask achieves the lowest Eval EER at 5.22%, while the Bands mask, effective in MaskedSpec, now shows increased Eval EER (8.78%). This suggests that masking at the feature representation level (MaskedFeature) introduces complexities different from those encountered in direct audio-level augmentation (MaskedSpec).
5.2 Pretrained Feature Extractor
The results presented in Table 5 emphasize the substantial influence of the pretrained feature extractor on the model’s performance. This is further evident in the fusion results, where models with different pretraining weights outperformed individual models, highlighting the complementary nature of diverse pretrained representations. Notably, while XLS-R 128 was trained on a larger and more linguistically diverse dataset, XLS-R 53 surprisingly achieved a lower Equal Error Rate (EER) during evaluation. This result suggests a more favorable compatibility of XLS-R 53 with the ASVSpoof5 evaluation dataset. In contrast, the Large-960h model, pretrained exclusively on the English-only Librispeech dataset, underperformed relative to the multilingual models. Given that the evaluation dataset is also English-only, these results suggest that multilingual pretraining improves the adaptability of the model and enhances the feature extractor’s ability to generalize.
Pretraining on the original ASVSpoof5 dataset, despite its task-specific focus on English and spoofed audio, surprisingly yielded a high EER. This outcome suggests a potential mismatch between the pretrained feature extractor and the requirements of the evaluation dataset. A plausible explanation is that the small size and limited diversity of the ASVSpoof5 dataset restrict the generalization capabilities of the pretrained model, which are critical for effective self-supervised learning.
When compression augmentations are introduced in the ASVSpoof5+ configuration, the evaluation EER improves significantly. This improvement suggests that the suboptimal performance observed with the ASVSpoof5 pretraining was likely due to the limited size and lack of variability in the dataset—factors critical for effective self-supervised learning. The introduction of compression augmentations enhances the model’s ability to generalize, thereby mitigating the challenges posed by the dataset’s constraints.
5.3 Mixed Augmentations
Mixed augmentations introduce variability that originally design to support generalization across diverse conditions, though excessive disruption can obscure key features and complicate learning. Table 3 illustrates that LPF and feature normalization generally improve Dev EER across mask types, but their impact on Eval EER is less consistent. For instance, Gauss masking with feature normalization achieves a low Dev EER (0.94%) but shows a modest Eval EER (5.82%), suggesting careful augmentations might support effective training metrics without assured generalization.
While feature normalization seems more beneficial with Gauss mask, LPF may better support simpler masks like Bands and Squares, likely by minimizing the disruption of essential features. These observations suggest that well-selected augmentations could enhance robustness, though the alignment between Dev and Eval metrics remains variable.
6 Conclusions and Future Work
This work introduces a deepfake speech detection model with a hybrid architecture, combining feature extractor architecture and a classifier head, trained end-to-end on the ASVSpoof5 dataset set. Despite limitations in dataset size, our model incorporates data augmentation and training strategies to achieve robust performance. Augmentation techniques applied at audio and feature levels have been shown to impact model performance in distinct ways, enhancing the variability within the training process. The MaskedSpec and MaskedFeature augmentation strategies, specifically utilizing the Bands and Gauss masks, not only enhance the model’s robustness but also significantly strengthen its performance in fusion configurations. These augmentations enable the model to generalize better across diverse conditions, contributing to improved overall effectiveness in handling complex synthetic audio scenarios. The results indicate that different mask shapes are well-suited to address various types of speech distortions.
Our findings indicate that multilingual pretraining, as utilized in the XLS-R 128 and XLS-R 53 models, enable our model to achieve lower evaluation EERs compared to monolingual models. Direct pretraining on ASVSpoof5 with the inclusion of compression augmentations (ASVSpoof5+) demonstrates strong performance, indicating that incorporating compression augmentations during pretraining is particularly effective when working with small, single-language datasets.
Notably, our model’s end-to-end training approach and augmentations establish a SoTA results for ASVSpoof5 benchmark in closed-set conditions and achieve even lower EERs when evaluated without closed conditions. Extension of this work can further explore pretraining augmentations and methods, as our findings suggest that different datasets capture diverse audio features. Extending augmentation strategies, such as incorporating masking augmentations directly into the pretraining phase, may further enhance the model’s robustness to distortions commonly introduced by spoofing techniques. Additionally, evaluating alternative classifier heads, with a focus on reducing model size while maintaining or improving performance, could lead to a more efficient model suitable for real-time applications.